OpenStudy (anonymous):

Elle is pushing a box of mass 7.0 kilograms with the force of 25 newtons. If the force of friction is 2.6 newtons, what is the value of acceleration of the box? A)0.7 meters/second B)2.6 meters/second C)3.2 meters/second D)3.6 meters/second

OpenStudy (***[isuru]***):

 Hi, resultant force = applied force - friction = 25N - 2.6N = 22.4N According to Newton's second law F = ma 22.4N = 7 Kg*a a = 22.4/7 = 3.2 m/s^2 So the choice is C